mbega15:

y−11x=x2−12 and y=12x+30

4evablood:

Are there answer choices by any chance?

isuckatenglish:

y=12x+30 do 1y-30 then do 12x divided by the answer you got

isuckatenglish:

idk i tried lol

isuckatenglish:

havent done that in a while

isuckatenglish:

then 11x - 2x

4evablood:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @isuckatenglish then 11x - 2x \(\color{#0cbb34}{\text{End of Quote}}\) wait til the person a reply

isuckatenglish:

okay ik the second one better tho

isuckatenglish:

y−11x=x2−12 +11+2=13x y=13x+-12 12+1+13y 13x=13y ANSWER IS 1

prettygirlgang:

give me 1 sec

prettygirlgang:

u can use this link to help u https://www.tiger-algebra.com

isuckatenglish:

ik the second one is right 1stone wrong

jhonyy9:

y -11x = x^2 -12 y = 12x +30 do you need solve the sistem for x and y ?

jhonyy9:

@supie